OpenStudy (anonymous):

if f(x)=9x(cubed) +9x(squared)-x+c and f(-3)=1. whats the value of C

OpenStudy (zehanz):

\[f(x)=x^3+9x^2-x+c\]\[f(-3)=(-3)^3+9(-3)^2--3+c=1\]Now do the calculations (without calculator, of course!) and solve for c!

OpenStudy (anonymous):

i get confused at the end.. because i got -27-81+3+c, right?

OpenStudy (anonymous):

-105+c?

OpenStudy (abb0t):

Integrate the function with respect to x!

OpenStudy (zehanz):

Sorry, typo :( \[f(x)=9x^3+9x^2-x+c\]\[f(-3)=9(-3)^3+9(-3)^2--3+c=1\] So it's 9*27+9*9+3+c=1 243+81+3+c=1 ... ...

OpenStudy (anonymous):

-326?

OpenStudy (zehanz):

yup

OpenStudy (anonymous):

Nope. c = 160 \[9(-3)^3+9(-3)^2-(-3)+c=1 \]\[9(-27)+9(9)+(3)+c=1 \]\[-243+81+3+c=1\]\[-243+84+c=1\]\[-159+c=1\]\[c=160\] Refer to the attachment for a solution using Mathematica.

OpenStudy (zehanz):

Mathematica seems overkill to me ;) I admit it had to be -243 instead of 243, so the final answer is 160. And then I had the nerve to say do not use a calculator... Well, if the numbers are more than 100 away from 0 then I'd better use the damn thing!

OpenStudy (anonymous):

Well look at this way: should NASA have hired 10,000 math slaves doing pencil and paper hand calculations to support the space shuttle flights in real time? Without pre-programmed computers, and fast ones at that, no space program.